How is the compression factor calculated? Show the derivation of the feasible equations for determining the compression factor that can be used.


1
Expert's answer
2021-01-27T04:35:28-0500

Compressibility factor, usually defined as Z = pV/RT, is unity for an ideal gas.

In most engineering work, the compressibility factor is used as a correction factor to ideal behavior. Thus, vreal = Z is used to calculate the actual volume, vreal, as the product of the compressibility factor and the ideal gas volume, all at the same pressure and temperature. Z is most commonly found from a generalized compressibility factor chart as a function of the reduced pressure, pr = p/pc, and the reduced temperature, Tτ = "\frac{T}{T_c}" where pc and Tr are the reduced variables and the subscript 'c' refers to the critical point.